40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 4 20:13:39 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors_1_Feb_2019_2230_PST.pdfETMY_UL.pdfETMY_sensors_timeDomain.pdf
    Reply  Thu Feb 7 19:34:18 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors.pdfpostDiag.pdf
       Reply  Fri Feb 8 20:48:52 2019, gautam, Update, LSC, IFO recovery IMG_7307.JPG
          Reply  Mon Feb 11 15:41:49 2019, gautam, Update, LSC, TRY 60 Hz solved - but clipping persists 
             Reply  Mon Feb 11 16:38:34 2019, gautam, Update, LSC, ETMY OL calibration updated OL_calib_ETMY_PERROR.pdfOL_calib_ETMY_YERROR.pdf
             Reply  Fri Feb 15 18:42:57 2019, gautam, Update, LSC, TRY 60 Hz solved  IMG_7323.JPG
Message ID: 14441     Entry time: Thu Feb 7 19:34:18 2019     In reply to: 14433     Reply to this: 14445
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY suspension oddness 

I did some tests of the electronics chain today.

  1. Drove a sine-wave using awggui to the UL-EXC channel, and monitored using an o'scope and a DB25 breakout board at J1 of the satellite box, with the flange cable disconnected - while driving 3000 cts amplitude signal, I saw a 2 Vpp signal on the scope, which is consistent with expectations.
  2. Checked resistances of the pin pairs corresponding to the OSEMs at the flange end using a breakout board - all 5 pairs read out ~16-17 ohms.
  3. Rana pointed out that the inductance is the unambiguous FoM here: all coils measured between 3.19 and 3.3 mH according to the LCR meter...

Hypothesising a bad connection between the sat box output J1 and the flange connection cable. Indeed, measuring the OSEM inductance from the DSUB end at the coil-driver board, the UL coil pins showed no inductance reading on the LCR meter, whereas the other 4 coils showed numbers between 3.2-3.3 mH. Suspecting the satellite box, I swapped it out for the spare (S/N 100). This seemed to do the trick, all 5 coil channels read out ~3.3 mH on the LCR meter when measured from the Coil driver board end. What's more, the damping behavior seemed more predictable - in fact, Rana found that all the loops were heavily overdamped. For our suspensions, I guess we want the damping to be critically damped - overdamping imparts excess displacement noise to the optic, while underdamping doesn't work either - in past elogs, I've seen a directive to aim for Q~5 for the pendulum resonances, so when someone does a systematic investigation of the suspensions, this will be something to look out for.. These flaky connectors are proving pretty troublesome, let's start testing out some prototype new Sat Boxes with a better connector solution - I think it's equally important to have a properly thought out monitoring connector scheme, so that we don't have to frequently plug-unplug connectors in the main electronics chain, which may lead to wear and tear.

The input and output matrices were reset to their "naive" values - unfortunately, two eigenmodes still seem to be degenerate to within 1 mHz, as can be seen from the below spectra (Attachment #1). Next step is to identify which modes these peaks actually correspond to, but if I can lock the arm cavities in a stable way and run the dither alignment, I may prioritize measurement of the loss. At least all the coils show the expected 1/f**2 response at the Oplev error point now. The coil output filter gains varied by ~ factor of 2 among the 4 coils, but after balancing the gains, they show identical responses in the Oplev - Attachment #2.

Attachment 1: ETMY_sensors.pdf  34 kB  Uploaded Fri Feb 8 02:17:26 2019  | Hide | Hide all
ETMY_sensors.pdf
Attachment 2: postDiag.pdf  52 kB  Uploaded Mon Feb 11 09:33:06 2019  | Hide | Hide all
postDiag.pdf
ELOG V3.1.3-